# TuneNSearch: a hybrid transfer learning and local search approach for solving vehicle routing problems

### v2

@@ 5 Experimental results
Baselines: To evaluate the performance of TuneNSearch, we compared it against four state-of-the-art baseline approaches: PyVRP, OR-Tools Routing library, MD-MTA and POMO. For PyVRP and OR-Tools, we set time limits of 300, 600 and 1800 seconds for problems of sizes $n=20,50,100$, respectively. Both methods were parallelized across 32 CPU cores, following prior research (Kool et al., 2019; Zhou et al., 2024a).
Training and hyper-parameters: We consider three MDVRP models trained on instances of sizes $n=20,50,100$, where the number of depots ($m$) was set to 2, 3, and 4, respectively. Each model was trained for 100 epochs, with one epoch consisting of 320000 training instances for $n\in{20,50}$, and 160000 for $n=100$. Batch sizes were set to 2000, 500 and 120 for each instance size, respectively. When fine-tuning our pre-trained MDVRP model, we consider 20 additional epochs for each VRP variant. For all baselines, we maintained the original hyper-parameters from their respective works. For TuneNSearch, we set the hidden dimension at $h_{x}=256$ and the hidden edge dimension at $h_{e}=32$. The number of encoder layers is $L=5$, and the fully connected layer in the encoder has a hidden dimension of 512. The number of heads in the MHA decoder is $H=16$, and the tanh clip size was set to 10, according to Bello et al. (2017). We employed the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $\eta=10^{-4}$. We provide a sensitivity analysis of key hyper-parameters in A.
@@ 5.1 Computational results on randomly generated instances
Table 1 provides a comparison between our pre-trained MDVRP model and MD-MTA, as well as between POMO (specialized for each specific VRP variant) and our fine-tuned models. The evaluation is based on 1280 randomly generated instances per problem and considers three key metrics: the average total distance traveled (Obj.), the average gap (Gap) compared to the best solutions found across all evaluated methods, and the computational time (Time) taken to solve all instances. The best results — i.e., the lowest average total distance traveled and gap (relative to PyVRP) — are highlighted in bold.
@@
Table 1: Experimental results on all VRP variants (* represents $0$ gap).
@@
Obj. Gap Time (m) Obj. Gap Time (m) Obj. Gap Time (m)
MDVRP PyVRP 4.494 * 200.134 7.489 * 401.243 11.345 * 1200.436
@@
For MDVRP instances, TuneNSearch consistently outperformed MD-MTA across all three problem sizes, even without employing the local search. When the local search is applied, the improvements became more pronounced, achieving minimal integrality gaps compared to PyVRP. Notably, even without x8 augmentation, the application of the local search significantly enhanced performance, allowing our model to perform better than OR-Tools and MD-MTA.
@@
We also conducted an additional set of experiments — using the same 100-node instances of Table 1 — where PyVRP and OR-Tools were limited to a similar computational time budget to that of TuneNSearch. Note that, in some variants, OR-Tools required more time than PyVRP and TuneNSearch, as this was the minimum time necessary to find a feasible solution. For TuneNSearch, we report results with the local search algorithm after inference. As shown in Table 2, when running under a comparable time budget, TuneNSearch outperformed PyVRP in 5 out of the 7 tested variants, achieving a lower average objective value overall. In contrast, OR-Tools did not prove to be competitive, performing significantly worse than both PyVRP and TuneNSearch in all variants. These results suggest that, in time-constrained applications, TuneNSearch is the most viable alternative.
@@
Obj. Gap Time (m)
MDVRP PyVRP 11.601 2.256 % 3.505
@@ 5.2 Computational results on benchmark instances
Table 3: Generalization on Cordeau’s benchmark instances (* represents $0$ gap).
@@
Avg. Gap 28.054 % 21.380 % 20.317 % 18.490 % 1.574 %
5.3 Ablation study
To evaluate the impact of the local search algorithm, we conducted an ablation study by varying the number of iterations performed after inference across all VRP variants, as outlined in Table 8. The results reveal that performance sensitivity to the number of iterations varies significantly between tasks. For example, tasks such as VRPTW, VRPB and OVRP demonstrated a higher sensitivity to changes in the number of iterations. In contrast, simpler variants like the CVRP and TSP exhibited much lower sensitivity. We believe this difference may be related to the level of constraints inherent to each task. More constrained tasks likely benefit more from additional iterations, as the search process requires more computational effort to explore the solution space effectively. Overall, we found that 50 iterations offer a good trade-off between solution quality and computational performance for most tasks. Beyond this point, further iterations lead to marginal gains, with the objective function improving at a much slower rate while imposing more computational time. Extending the local search for too many iterations would diminish the computational efficiency enabled by the reinforcement learning component, which goes against the intended purpose of our method.
Table 8: Ablation study on the number of local search iterations on 1280 randomly generated instances.
Number of iterations Problem $n$ = 20 $n$ = 50 $n$= 100
Obj. Time (m) Obj. Time (m) Obj. Time (m)
5 MDVRP 4.498 0.124 7.616 0.282 11.669 0.929
CVRP 4.969 0.129 9.481 0.252 16.368 0.793
VRPB 4.599 0.119 8.460 0.257 14.304 0.767
VRPL 4.991 0.123 9.481 0.259 16.378 0.828
OVRP 3.485 0.124 6.236 0.239 10.251 0.791
VRPTW 7.684 0.134 14.910 0.310 25.683 0.975
TSP 3.831 0.119 5.713 0.211 7.833 0.599
Average 4.865 0.125 8.842 0.259 14.641 0.812
10 MDVRP 4.493 0.143 7.594 0.365 11.646 1.107
CVRP 4.965 0.144 9.465 0.331 16.357 1.055
VRPB 4.596 0.126 8.436 0.328 14.288 1.097
VRPL 4.987 0.147 9.464 0.332 16.366 1.063
OVRP 3.482 0.149 6.215 0.318 10.217 0.999
VRPTW 7.654 0.148 14.792 0.427 25.518 1.296
TSP 3.831 0.129 5.712 0.253 7.830 0.784
Average 4.858 0.141 8.811 0.336 14.603 1.057
25 MDVRP 4.487 0.191 7.563 0.584 11.606 1.793
CVRP 4.963 0.153 9.441 0.555 16.332 1.763
VRPB 4.593 0.155 8.402 0.559 14.249 1.847
VRPL 4.984 0.162 9.443 0.570 16.338 1.804
OVRP 3.479 0.143 6.188 0.514 10.167 1.657
VRPTW 7.624 0.181 14.667 0.763 25.320 2.286
TSP 3.830 0.153 5.708 0.413 7.823 1.302
Average 4.851 0.163 8.773 0.565 14.548 1.779
50 MDVRP 4.486 0.251 7.541 0.938 11.570 2.902
CVRP 4.962 0.214 9.425 0.936 16.306 2.975
VRPB 4.592 0.228 8.382 0.963 14.212 2.945
VRPL 4.983 0.222 9.427 0.994 14.313 3.048
OVRP 3.478 0.200 6.173 0.866 10.125 2.746
VRPTW 7.610 0.274 14.592 1.394 25.178 4.027
TSP 3.830 0.168 5.705 0.676 7.816 2.190
Average 4.849 0.222 8.749 0.967 14.217 2.976
100 MDVRP 4.485 0.371 7.525 1.652 11.526 5.128
CVRP 4.962 0.349 9.410 1.711 16.278 5.396
VRPB 4.592 0.347 8.365 1.705 14.170 5.446
VRPL 4.983 0.354 9.412 1.771 16.282 5.460
OVRP 3.478 0.326 6.162 1.549 10.084 4.917
VRPTW 7.604 0.463 14.539 2.622 25.026 7.258
TSP 3.830 0.249 5.704 1.229 7.808 4.079
Average 4.848 0.351 8.731 1.748 14.453 5.383
@@ 5.4 Why pre-train on the MDVRP?
We argue that one of the main contributions of this paper is that pre-training the model on the MDVRP can lead to a better generalization across different single- and multi-depot tasks. To assess the effectiveness of pre-training the model on MDVRP data, we compared TuneNSearch with a model pre-trained on the CVRP. We evaluate both models on their zero-shot generalization performance to both single-depot and multi-depot variants. For the multi-depot variants, we incorporated the constraints outlined in Section 3.3 into the classic MDVRP, resulting in the following variants: MDVRP with backhauls (MDVRPB), MDVRP with duration limits (MDVRPL), MDVRP with open routes (MDOVRP), and MDVRP with time windows (MDVRPTW).
Like TuneNSearch, we also pre-trained the CVRP model for 100 epochs, using the same deep neural network architecture. Table 9 presents the inference results for both models in terms of the objective value and the performance gap between TuneNSearch and the CVRP pre-trained model. The evaluation was conducted on 1280 randomly generated instances for each VRP variant. In both cases, inference was performed greedily with instance augmentation. In these experiments, we did not apply the local search algorithm.
Results show that TuneNSearch performs just as well as the CVRP pre-trained model on single-depot variants, with a negligible performance gap between both methods. However, on multi-depot variants, TuneNSearch performs significantly better on all instance sizes. The performance gap becomes especially more pronounced as problem size increases: while the CVRP pre-trained model performs reasonably well on small instances (with 20 nodes), its performance degrades sharply on larger problems.
@@
Obj. Performance gap Obj. Performance gap Obj. Performance gap
@@
5.5 Impact of integrating the residual E-GAT with POMO
@@
Another key contribution of our work is the integration of POMO (Kwon et al., 2020) with the residual E-GAT encoder (Lei et al., 2022). The residual E-GAT, originally built on top of the attention model (Kool et al., 2019), improves the learning process by incorporating edge information and adding residual connections between layers. This allows the model to capture richer contextual information about the problem, resulting in more accurate attention coefficient computations. When combined with POMO, this improved encoding can lead to a better utilization of POMO’s multiple starting nodes, generating higher quality solutions. However, the impact of the residual E-GAT has not previously been evaluated in conjunction with POMO — a framework that has already demonstrated superior performance compared to the original attention model.
@@
In this subsection, we compare the performance of TuneNSearch with both the original POMO and residual E-GAT models. Our aim is to analyze how effective the encoding capability of each approach is, and how well they capture the graph features of the VRP. We begin by examining the training patterns of all three approaches, analyzing the evolution of the average objective function at each epoch, illustrated in Fig. 5. The models were trained in MDVRP instances with 20, 50 and 100 customers and 2, 3 and 4 depots, respectively. Across all three problem sizes, TuneNSearch exhibited a more efficient convergence. In comparison to the residual E-GAT model, the gap was a lot more noticeable, since it does not incorporate the exploitation of solutions symmetries and multiple starting nodes introduced by POMO.
@@
Table 10: Average inference results on 1280 randomly generated MDVRP instances: impact of integrating the residual E-GAT with POMO.

### v3

@@ 5 Experimental results
Baselines: To evaluate the performance of TuneNSearch, we compared it against four state-of-the-art baseline approaches: PyVRP, an extension of the state-of-the-art hybrid genetic search (HGS) (Vidal, 2022), OR-Tools Routing library, MD-MTA and POMO. For PyVRP and OR-Tools, we set time limits of 300, 600 and 1800 seconds for problems of sizes $n=20,50,100$, respectively. Both methods were parallelized across 32 CPU cores, following prior research (Kool et al., 2019; Zhou et al., 2024a).
Training and hyper-parameters: We consider three MDVRP models trained on instances of sizes $n=20,50,100$, where the number of depots ($m$) was set to 2, 3, and 4, respectively. Each model was trained for 100 epochs, with one epoch consisting of 320000 training instances for $n\in{20,50}$, and 160000 for $n=100$. The batch size $B$ was chosen based on GPU memory limits to make training more efficient and faster. For smaller problems (with $n=20$), we used $B=2000$, since the instances were simpler and allowed a more effective GPU utilization. For larger problems, with 50 and 100 nodes, we reduced $B$ to 500 and 120, respectively, since these instances required much more memory than the 20-node ones. When fine-tuning our pre-trained MDVRP model, we consider 20 additional epochs for each VRP variant. For all baselines, we maintained the original hyper-parameters from their respective works. For TuneNSearch, we set the hidden dimension at $h_{x}=256$ and the hidden edge dimension at $h_{e}=32$. The number of encoder layers is $L=5$, and the feed-forward layer in the encoder has a hidden dimension of 512. The number of heads in the MHA decoder is $H=16$, and the tanh clip size was set to 10, according to Bello et al. (2017). We employed the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $\eta=10^{-4}$. Across all experiments, we set the number of starting nodes $N$ to g1g-1. Essentially, this means that for every instance solved, g1g-1 different solutions are generated following the procedure explained in section 4.2. This choice follows a convention in the literature, from which most studies adopt the same value (Kwon et al., 2020; Lin et al., 2024; Li et al., 2024). It allows a broad exploration of the solution space, while keeping the total training time manageable. Finally, a sensitivity analysis of key hyper-parameters is provided in A.
@@ 5.1 Computational results on randomly generated instances
Table 1 provides a comparison between our pre-trained MDVRP model and MD-MTA, as well as between POMO (specialized for each specific VRP variant) and our fine-tuned models. The evaluation is based on datasets of 1280 randomly generated instances for each variant and instance size pair, and considers three key metrics: the average total distance traveled (Obj.), the average relative percentage deviation (RPD) compared to the best solutions found across all evaluated methods, and the computational time (Time) taken to solve all instances. The RPD for each method is computed as the average, over all instances in a dataset $D$, of the percentage deviation from the best-performing method on each instance (i.e., the method achieving the lowest objective on that instance):
$RPD=\frac{1}{|D|}\sum_{i=1}^{|D|}\frac{z_{i}-z_{best,i}}{z_{best,i}}\times 100\%$ (15)
where $z_{i}$ is the objective value of the evaluated method on instance $i$, and $z_{best,i}$ is the smallest objective value among all evaluated methods on instance $i$. The best results, indicated in bold, correspond to the lowest Obj. values and the lowest RPD (relative to HGS-PyVRP).
@@
Table 1: Experimental results on all VRP variants (* represents 0.000 % RPD).
@@
Obj. RPD Time (m) Obj. RPD Time (m) Obj. RPD Time (m)
MDVRP HGS-PyVRP 4.494 * 200.134 7.489 * 401.243 11.345 * 1200.436
@@
For MDVRP instances, TuneNSearch consistently outperformed MD-MTA across all three problem sizes, even without employing the local search. When the local search is applied, the improvements became more pronounced, achieving minimal RPDs compared to PyVRP. Notably, even without x8 augmentation, the application of the local search significantly enhanced performance, allowing our model to perform better than OR-Tools and MD-MTA.
@@
We also conducted an additional set of experiments using the same 100-node instances of Table 1, where PyVRP and OR-Tools were limited to a similar computational time budget to that of TuneNSearch. Note that, in some variants, OR-Tools required more time than PyVRP and TuneNSearch, as this was the minimum time necessary to find a feasible solution. For TuneNSearch, we report results with the local search algorithm after inference. As shown in Table 2, when running under a comparable time budget, TuneNSearch outperformed PyVRP in 5 out of the 7 tested variants, achieving a lower average objective value overall. In contrast, OR-Tools did not prove to be competitive, performing significantly worse than both PyVRP and TuneNSearch in all variants. These results suggest that, in time-constrained applications, TuneNSearch is the most viable alternative.
@@
Obj. RPD Time (m)
MDVRP HGS-PyVRP 11.601 2.256 % 3.505
@@ 5.2 Computational results on benchmark instances
Table 3: Generalization on Cordeau’s benchmark instances (* represents 0.000 % RPD).
@@
Avg. RPD 28.054 % 21.380 % 20.317 % 18.490 % 1.574 %
@@ 5.3 Why pre-train on the MDVRP?
We argue that one of the main contributions of this paper is that pre-training the model on the MDVRP can lead to a better generalization across different single- and multi-depot tasks. To assess the effectiveness of this approach, we compared TuneNSearch with a model pre-trained on the CVRP. We evaluate both models on their zero-shot generalization performance to both single-depot and multi-depot variants. Zero-shot generalization refers to the model’s ability to perform well on new VRP variants it has not seen during training, without any additional fine-tuning. For the multi-depot variants, we incorporated the constraints outlined in Section 3.3 into the classic MDVRP, resulting in the following variants: MDVRP with backhauls (MDVRPB), MDVRP with duration limits (MDVRPL), MDVRP with open routes (MDOVRP), and MDVRP with time windows (MDVRPTW).
Like TuneNSearch, the CVRP-based model was trained for 100 epochs, using the same deep neural network architecture. Table 8 presents the inference results for both models in terms of the objective value and the performance deviation of TuneNSearch relative to the CVRP model. The evaluation was conducted on 1280 randomly generated instances for each VRP variant. In both cases, inference was performed greedily with instance augmentation. In these experiments, we did not apply the local search algorithm.
Results show that TuneNSearch performs similarly to the CVRP-based model on single-depot variants, with a negligible performance deviation between both methods. However, on multi-depot variants, TuneNSearch performs significantly better on all instance sizes. The performance deviation becomes especially more pronounced as problem size increases: while the model pre-trained on the CVRP performs reasonably well on small instances (with 20 nodes), its performance degrades sharply on larger problems.
@@
Obj. Performance deviation Obj. Performance deviation Obj. Performance deviation
@@
5.4 Local search effectiveness assessment
To evaluate the impact of the local search algorithm, we conducted a sensitivity analysis on its convergence behavior over successive iterations, shown in Fig. 5. This analysis allowed us to examine how solution quality evolves with increasing iterations and how quickly the algorithm converges. Specifically, we tracked the evolution of the best solution found by our method over 400 iterations of the local search for all tested problem variants, on instances with 20, 50 and 100 nodes. For these experiments, we used the same models as in Section 5.1 (with label Ours (ls.) in Table 1). For each variant-size combination, the analysis included 1280 randomly generated instances, and the plots report the average of the best solutions obtained by the model up to each iteration. To further understand the convergence behavior of our method, we performed an ABC analysis for each variant-size pair, represented by the green, yellow, and red shaded areas in each subplot of Fig. 5. The ”A” group (green) shows the early iterations, during which the algorithm achieves majority of its improvement (specifically, 70% of the total improvement relative to the final solutions at iteration 400). The ”B” group (yellow) covers the iterations during which the algorithm reaches a 90% of the final improvement, while the ”C” group (red) represents the final 10% of improvement.
Overall, we can identify different patterns through this sensitivity analysis. On smaller problems, across all variants tested, in most cases the algorithm reached the ”B” group before the 10th iteration, showing a very fast convergence. For problems with 50 nodes, this happened, in general, between iterations 20 and 30, while for problems with 100 nodes, about 70 iterations were needed, in the worst case. We can also notice that some variants displayed a much smoother convergence than others. For example, tasks such as VRPTW and VRPL demonstrated a higher sensitivity to changes in the number of iterations. In contrast, simpler variants like the TSP exhibited much lower sensitivity, with a very rapid convergence even on 100-node instances.
Overall, we found that 50 iterations offer a good trade-off between solution quality and computational performance. Beyond this point, further iterations yield marginal improvements, with the objective function improving at a much slower rate while imposing more computational time. Since TuneNSearch was designed intended for large-scale, time-sensitive applications where scalability is critical, extending the local search for too many iterations would undermine the efficiency benefits introduced by the reinforcement learning component, which goes against our primary design goal.
Figure 5: Local search convergence of all variants across 1280 randomly generated instances.
Having analyzed the convergence behavior of the local search algorithm, we now turn our attention to the role of solution initialization. In particular, we examine whether better initial solutions significantly affect the final solution quality. TuneNSearch benefits from a neural-based model to generate high-quality solutions, starting from more promising regions of the search space and accelerating convergence. Therefore, to assess the effectiveness of this approach, we compare it with two other solution initialization strategies. The first initialization strategy (Random) begins with an empty solution, and constructs routes sequentially. Customer nodes are assigned to the current route in a random order, as long as the vehicle has enough capacity to meet their demand. Once a route can no longer accommodate additional customers, a new route is started, and the process continues until all customers are assigned. The resulting solutions are then improved through the same local search procedure as in our method. In the second initialization strategy (Greedy), routes are also built sequentially. However, instead of assigning customers randomly, each customer is chosen based on the nearest neighbor method, that is, the closest unvisited customer is chosen, provided the vehicle has enough capacity.
As in previous experiments, each approach was tested on 1280 randomly generated instances for the different variants solved in Section 5.1. The experiments included problems with 20, 50, 100 and 200 nodes. We evaluated the performance of each approach using three different metrics: Obj., the average objective function across all 1280 instances; Win rate %, the percentage of instances in which each of the approaches outperformed the others; Time (m), the total computational time (in minutes) required to solve all 1280 instances. Note that the Win rate % values do not always sum to 100% since all three approaches may achieve identical solutions on the same instance. All results are shown in Table 9.
Overall, our neural-based initialization consistently outperformed the other two strategies across almost all instance sizes and variants. We can also notice that the effectiveness of TuneNSearch becomes increasingly pronounced as instance size grows. For smaller problems, all strategies perform similarly because the search space is limited, allowing the local search to refine even lower-quality initial solutions effectively. However, on problems with 100 and 200 nodes, our method outperformed the other in the majority of instances, achieving a Win (%) of around 70-80% compared to the other initialization strategies. This suggests that the neural-based initialization allows the local search algorithm to significantly bypass the initial effort required to get close to promising regions of the solution space.
Table 9: Comparison of neural-based, random and greedy solution initialization strategies on 1280 randomly generated instances.
Method $n=20$ $n=50$ $n=100$ $n=200$
Obj. Win (%) Time (m) Obj. Win (%) Time Obj. Win % Time Obj. Win (%) Time
MDVRP Ours 4.504 4.14% 0.271 7.525 34.69% 1.059 11.485 75.08% 2.989 19.567 86.17% 11.381
Random 4.508 0.94% 0.266 7.555 14.30% 1.237 11.640 12.19% 3.388 19.910 5.55% 11.492
Greedy 4.507 1.02% 0.251 7.552 15.55% 1.249 11.635 12.73% 3.412 19.887 8.28% 11.648
CVRP Ours 4.978 1.88% 0.251 9.398 35.23% 1.024 16.235 84.77% 2.926 29.509 88.36% 11.385
Random 5.011 0.08% 0.248 9.419 14.77% 1.274 16.418 8.83% 3.563 29.871 5.23% 12.231
Greedy 4.989 0.62% 0.251 9.419 15.70% 1.281 16.425 5.94% 3.591 29.861 6.09% 12.468
VRPB Ours 4.606 0.55% 0.257 8.383 33.83% 1.012 14.176 71.64% 2.898 25.209 77.89% 10.884
Random 4.606 0.23% 0.248 8.400 20.08% 1.262 14.305 13.44% 3.684 25.467 9.14% 11.901
Greedy 4.606 0.39% 0.248 8.403 17.19% 1.274 14.311 13.36% 3.514 25.471 9.22% 11.992
VRPL Ours 5.000 1.80% 0.257 9.396 37.50% 1.046 16.240 82.73% 3.052 29.504 90.62% 12.194
Random 5.002 0.31% 0.242 9.422 15.78% 1.262 16.422 7.97% 3.551 29.877 4.61% 12.338
Greedy 5.002 0.39% 0.245 9.420 14.14% 1.274 16.420 8.67% 3.591 29.880 4.45% 12.463
OVRP Ours 3.484 0.39% 0.245 6.160 19.84% 0.950 10.105 46.41% 2.761 17.555 51.41% 10.926
Random 3.484 0.08% 0.230 6.166 13.20% 1.116 10.139 26.72% 3.198 17.622 22.73% 11.033
Greedy 3.484 0.16% 0.236 6.166 15.39% 1.129 10.137 24.30% 3.238 17.618 24.22% 11.245
VRPTW Ours 7.657 7.27% 0.295 14.728 28.12% 1.439 25.174 36.88% 3.974 44.946 37.03% 14.438
Random 7.662 1.56% 0.319 14.746 20.08% 1.892 25.198 26.48% 4.922 44.926 28.12% 15.014
Greedy 7.664 1.41% 0.322 14.745 20.47% 1.904 25.198 28.20% 4.945 44.911 30.86% 15.308
TSP Ours 3.824 0.00% 0.233 5.688 28.59% 0.809 7.780 96.64% 2.379 10.900 99.61% 9.232
Random 3.824 0.00% 0.207 5.716 1.09% 0.947 8.019 1.25% 2.780 11.528 0.00% 9.820
Greedy 3.824 0.00% 0.204 5.715 2.19% 0.951 8.003 1.48% 2.769 11.395 0.39% 9.804
Average Ours 4.865 2.29% 0.260 8.754 31.11% 1.050 14.456 70.59% 2.997 25.313 75.87% 11.481
Random 4.871 0.46% 0.251 8.775 14.19% 1.281 14.592 13.84% 3.582 25.600 10.77% 11.988
Greedy 4.868 0.57% 0.251 8.774 14.38% 1.294 14.590 13.53% 3.578 25.575 11.93% 12.121
5.5 Impact of integrating the E-GAT with POMO
@@
Another key contribution of our work is the integration of POMO (Kwon et al., 2020) with the E-GAT encoder (Lei et al., 2022). The residual E-GAT, originally built on top of the attention model (Kool et al., 2019), aims at improving the learning process by incorporating edge information of all nodes in a problem. This allows the model to capture richer contextual information, resulting in more accurate attention coefficient computations. When combined with POMO, this improved encoding can lead to a better utilization of POMO’s multiple starting nodes, generating higher quality solutions.
@@
In this subsection, we compare the performance of TuneNSearch with both the original POMO and residual E-GAT models. Our aim is to analyze how effective the encoding capability of each approach is, and how well they capture the graph features of the VRP. We begin by examining the training patterns of all three approaches, analyzing the evolution of the average objective function at each epoch, illustrated in Fig. 6. The models were trained in MDVRP instances with 20, 50 and 100 customers and 2, 3 and 4 depots, respectively. Across all three problem sizes, TuneNSearch exhibited a more efficient convergence. In comparison to the residual E-GAT model, the difference was a lot more noticeable, since it does not incorporate the exploitation of solutions symmetries and multiple starting nodes introduced by POMO.
@@
Table 10: Average inference results on 1280 randomly generated MDVRP instances: impact of integrating E-GAT with POMO.